### v1

# Multi-neutron correlations in light nuclei via ab-initio lattice simulations

Abstract
The quest to understand multi-neutron systems has a long history, and recent experimental efforts aim to probe candidate four-neutron configurations in neutron-rich light nuclei such as 8He and 7H via quasi-free knockout reactions. However, the ground-state energies of the hydrogen isotopes 6H and 7H are not yet well constrained, with substantial discrepancies across experimental analyses and theoretical predictions. Using ab initio nuclear lattice effective field theory with an ensemble of 282 chiral two- and three-nucleon forces, we perform an uncertainty-quantified analysis of the ground-state energies of 6H and 7H. The marginal posteriors suggest single-neutron separation energy MeV, disfavoring sequential decay and pointing towards direct emission. Intrinsic densities indicate triton- and -like clusters in 7H and 8He, respectively. By computing two-body and reduced four-body correlation functions, we find that the valence neutrons in the surface region of these systems form compact dineutrons that predominantly organize into symmetric dineutron-dineutron configurations, with only a small but non-negligible fraction assembling into more compact tetraneutron-like substructures. In 7H, these components account for roughly 95% and 5% of the sampled four-neutron configurations, respectively, and 8He exhibits a similar hierarchy. For these configurations, we also extract the corresponding spatial and angular correlation patterns among the nucleons. These results provide nuclear-structure insights into the debate surrounding four-neutron clusters and complement ongoing experimental searches for tetraneutron signatures in light nuclei.
today
Introduction: The nature of multi-neutron systems, in particular candidate three-neutron () and four-neutron () clusters, has been the subject of long-standing interest and controversy for more than half a century [1]. Experimentally, our knowledge of such systems relies predominantly on indirect kinematic reconstructions in breakup [2], double charge-exchange [3], and quasi-free knockout reactions [4], which are often discussed within a picture invoking preformed four-neutron configurations in neutron-rich projectiles. A fully exclusive observation of a correlated multi-neutron cluster remains a major challenge and the focus of ongoing efforts [5]. Such systems provide a unique window into many-body correlations in neutron-rich nuclei and dilute nuclear matter, and their direct exploration would clarify both their existence and their internal multi-neutron correlations, thereby linking multi-neutron correlations to more general many-body correlations in nuclei and nuclear matter.
More generally, correlations among nucleons in atomic nuclei are central to our understanding of nuclear structure and reactions, as well as nuclear astrophysics [6, 7]. Over the past decades, major experimental progress has been made in probing these correlations, including the detailed characterization of short-range correlated (SRCs) pairs [8, 9], long-range pairing correlations, manifest in halo nuclei and in nuclear superfluidity [10, 11, 12, 13, 14, 15, 16, 17, 18, 20, 19, 21], observations of correlated two-nucleon emission from exotic nuclei [22, 24, 25, 26, 23], and signatures of multi-nucleon correlations in finite nuclei [28, 27]. At the same time, modern radioactive ion beam facilities have produced multi-proton [23] and multi-neutron [29, 30, 31, 32] emitters, yet direct and exclusive observations of correlated multi-neutron emission remain severely limited by detector efficiency and acceptance, posing a major experimental challenge for future experiments and highlighting the need for complementary theoretical approaches.
On the theoretical side, two-body correlations have been studied extensively, including their spatial structure [33, 34, 35, 36, 37, 38, 39, 40, 41] and dynamical properties [42, 43, 44, 45, 46]. Multi-nucleon correlations beyond the two-body level have primarily been investigated within Hartree–Fock–Bogoliubov approaches [47], few-body models [48], antisymmetrized molecular dynamics (AMD) calculations with explicit cluster projection [49, 50], and microscopic cluster calculations based on the generator coordinate method [51, 52, 53, 54], all of which have yielded important insights into dineutron () and – correlations in neutron-rich nuclei. In the SRCs regime, recent ab initio quantum Monte Carlo studies have probed three-nucleon correlations in the high-momentum components of nuclear wave functions [55]. However, a direct calculation of four-body (and higher) correlation functions in nuclear systems remains highly nontrivial, because reduced density matrices beyond the two-body level suffer from a combinatorial explosion in dimensionality and memory requirements. Further progress may thus be achieved with methods that can access -body correlations without explicit construction of the full -body density matrix. In this work, we exploit auxiliary-field Monte Carlo (AFMC) sampling to compute multi-neutron observables via Euclidean-time projection.
Motivated by quasi-free knockout experiments [56, 57, 58], which have revealed evidence for preformed clusters in neutron-rich nuclei [27] and have suggested that quasi-free knockout can access four-neutron configurations that closely resemble the intrinsic subsystem of the projectile nucleus [4], we focus on neutron-rich 7H and 8He. These cluster-core systems with four valence neutrons provide unique laboratories for studying emergent correlations, and our analysis of their internal correlations offers complementary nuclear-structure insights into the long-standing experimental and theoretical controversy surrounding and systems (see Ref. [1] and references therein and see Refs. [59, 4, 61, 60, 62] for recent developments). Here, we present an ab initio investigation of the neutron-rich nuclei 7H and 8He using nuclear lattice effective field theory (NLEFT) [63], with full -body correlations. We identify compact surface dineutrons that organize into two representative four-neutron configurations and quantify the geometry and probability of these configurations from two- and four-neutron correlation functions.
Methods: NLEFT provides a powerful framework to capture the full set of many-body correlations through AFMC simulations, but its application to high-fidelity chiral interactions has been hindered by severe sign problems [63, 64, 65]. With recent advancements via the wavefunction matching method, chiral forces up-to-and-including N3LO can now be treated largely free from sign problems [66]. In this work, we employ a set of N3LO chiral interactions that include both nucleon–nucleon (NN) and three-nucleon (3N) forces. These interactions are obtained through history matching [67, 68, 69, 66] that systematically incorporates uncertainties from NN scattering data, the truncation of the chiral hierarchy, and the binding energies of selected nuclei. The main results are derived from calculations using the best-fit interaction, while the corresponding uncertainties and correlations among observables are quantified using Bayesian posterior predictive distributions (PPD) constructed from an ensemble of 282 non-implausible chiral interactions, as detailed in [70].
To probe observables associated with nuclear density, particularly those related to intrinsic structure and correlation patterns, we employ the pinhole method [71]. This approach enables sampling of the spatial coordinates, spins, and isospins of all nucleons according to the quantum amplitudes of the -body density within the auxiliary-field framework. With these pinhole configurations, a wide range of observables of physical interest can be evaluated, as described in [70]. In the following, expectation values are denoted by . In order to reveal the intrinsic structure of the nucleus, an appropriate alignment scheme for the intrinsic density is often employed, as in Refs. [72, 71, 73, 74, 75]. Here, we identify candidate clusters (such as an particle or a triton) by selecting the most probable grouping of nucleons that matches the cluster configuration within each sampled pinhole configuration. Once the clusters are identified, we realign the density profile with respect to the cluster positions, thereby making the intrinsic spatial structure of the nucleus more apparent. Further technical details are provided in [70].
Following the identification of the intrinsic clusters, we extend our analysis to multi-neutron correlations. In contrast to compact clusters, these correlations exhibit more delocalized patterns. We begin with the two-body correlation function,
where denote the distances of the two nucleons from the center of mass (c.m.), is the opening angle between them, and represents the angle between the relative position vector of the two nucleons and the bisector of their opening angle. The first form, similar to those used in previous studies [76, 34, 36], characterizes the spatial distribution of two nucleons in terms of their relative distances from the c.m. and opening angle. The appropriate Jacobian is included when transforming from the pinhole coordinates. The second form, although it lacks an explicit radial-distance dependence, highlights the radial asymmetry between the two nucleons. When , the configuration is approximately symmetric, corresponding to two nucleons at similar distances from the c.m. This quantity is further utilized in the subsequent analysis of multi-neutron correlations.
In a four-body subsystem, correlations span many degrees of freedom, including individual coordinates and all pairwise opening angles, making it difficult to extract clear structural information directly. Building on the two-body correlations and symmetry considerations, we construct reduced projections that capture the essential features of the underlying multi-nucleon correlations. In particular, we define a reduced four-body correlation function
where label the nucleons, the sum runs over pairs of disjoint nucleon pairs and , and the angles and for each pair are defined in the same way as in the two-body case. The angle is the opening angle between the c.m. vectors of the two pairs, while denotes the torsion angle between the relative position vectors of nucleons in each pair. This definition projects the four-body density onto configurations in which two disjoint nucleon pairs exhibit a specified angular geometry, as schematically illustrated in [70].
In practice, we perform AFMC simulations on a periodic cubic lattice with side length of and spatial lattice spacing of fm, corresponding to a momentum cutoff of MeV. The temporal lattice spacing is set to MeV-1 for the Euclidean-time evolution. The initial states are chosen as the product of the single-particle harmonic oscillator (HO) wavefunctions. For the hydrogen isotopes, the HO frequency of the valence neutrons is reduced to generate more diffuse wavefunctions and to accelerate convergence. Calculated energies are extrapolated to the infinite Euclidean-time limit using the formalism outlined in [70]. Finite-volume effects have been explicitly examined, and their impact is found to be well within the statistical uncertainties [70]. In the pinhole calculation, the point-nucleon coordinates are convolved with a Gaussian [71] of width fm to account for the finite size of the nucleons [85]. Each pinhole configuration undergoes 3000 Gaussian smearing iterations to enhance the smoothness and reduce lattice artifacts in the density profiles. Unless stated otherwise, density-related observables are evaluated at a Euclidean projection time of MeV-1, ensuring a balance between convergence and statistical precision, following Refs. [71, 74, 75, 86].
Results:
The calculated energies of 6,7H and 8He are shown in Fig. 1 and compared with experimental data. For 8He, multichannel calculations projected onto the irrep of the cubic group [87, 88] yield energies of MeV and MeV for the and states, respectively, consistent with the experimental values MeV [89] and MeV [90]. In addition, the lattice simulations reveal a dineutron correlation in 8He that emerges and persists at sufficiently long Euclidean times, indicative of configuration mixing beyond the initial pure filling. A detailed discussion of this pattern is given below. This aligns with reaction [91, 92, 93] and structure [94, 53, 54] studies that indicate configuration mixing and deformation in 8He, and with the newly observed cluster-like excitation, for which strong – components have been discussed [90]. Extending to hydrogen isotopes, we find a ground state (g.s.) in 6H, indicating occupancy and a breaking of the subshell closure, which is conducive to dineutron correlations [36]. Experimentally, the reported 6,7H g.s. energies and widths vary across reactions (see Table S2 in [70]), and two advanced theoretical approaches predict 6,7H g.s. energies that differ by about 7 MeV [95, 96], underscoring the sensitivity to the interaction and many-body treatment. Incorporating full -body correlations, we perform a Bayesian analysis over 282 non-implausible chiral interactions to quantify uncertainties. The resulting g.s. energy medians and credible intervals for 6,7H are MeV and MeV, respectively, within the experimental ranges. A quantitative treatment of resonance widths requires a dedicated analysis and is beyond the present scope. The marginal posterior for the one-neutron separation energy of 7H suggests a positive at the 1 level, MeV, which kinematically disfavors sequential decay via 6H and thereby makes direct multi-neutron emission channels comparatively more relevant. These findings, together with experimental and theoretical indications [4, 61] of dineutron clustering in dynamics, underscore the emergence of enhanced multi-neutron correlations in 7H and 8He, from dineutron to possible tetraneutron-like four-body configurations, and set the stage for the correlation analysis that follows.
Fig. 2 presents the intrinsic nucleon densities and two-neutron () correlations in 7H, taken as a representative case of multi-neutron dynamics near the dripline (see [70] for benchmarks in 6He and complementary analysis of 8He). The intrinsic frame is defined with the triton c.m. located on the negative axis, so that the valence neutrons in this convention predominantly appear in the upper (positive-) half of the panels. The nucleon and proton density maps on the and slices reveal a compact triton cluster surrounded by four diffuse valence neutrons. This clustered geometry is consistent with a core–plus–valence-neutron picture as commonly employed in cluster and many-body approaches. The interpretation is further supported by Fig. 2(b), which displays pair correlations for all six neutrons relative to the triton c.m. and reveals two distinct components. The inner component is concentrated at short distances with only mild angular modulation, analogous to the correlations in 3H [70].
The outer component emerges at larger radii, where pronounced enhancements are observed at both small and large opening angles between the two neutrons. These two distinct components indicate that our core–valence partition has only a minor impact on the subsequent correlation analysis and, in particular, that it allows us to analyze the correlations among the valence neutrons. Motivated by this clear separation, we extract the two-body correlations among the valence neutrons in the 7H c.m. frame, Fig. 2(c), as commonly used in experimental analyses, and resolve more detailed features of the valence-neutron angular correlations. The resulting bimodal angular pattern closely resembles that obtained in Fig. 2(b), indicating the robustness of the dineutron correlations, and shows pronounced maxima around and at fm. At small , the enhancement indicates the presence of compact dineutron-like pairs in the surface region. We further extract the typical size for pairs with small , as detailed in Refs. [70]. The comparable strength at large contrasts with systems such as 6He, where only a relatively weak large- component has been reported [34, 36, 97]. For 7H, each pinhole configuration contains several neutron pairs, so the two-body correlation function represents a statistical distribution over all pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. The enhanced large- component indicates that, besides compact surface dineutrons, neutron pairs also populate more extended, back-to-back geometries. This pattern is qualitatively consistent with configurations in which the four valence neutrons form two dineutron-like clusters located on opposite sides of the core, in line with a more symmetric – halo geometry inferred from experimental analyses of 8He [98] and supported by microscopic calculations of 8He and 7H that exhibit – clustering [47, 48, 49, 50, 51, 52, 53, 54]. Such bimodal angular correlations are strongly suggestive of – clustering and may represent a microscopic mechanism underlying enhanced correlations in 7H and 8He.
The structure of configurations has been explored analytically using simple schematic models [99, 100] and in intuitive cluster descriptions that treat the four neutrons as two correlated dineutron clusters [47, 48, 49, 50, 51, 52, 53, 54]. However, a systematic characterization of four-neutron geometries at the nucleonic level remains limited, and an explicit four-body treatment is required to access detailed spatial and angular information. In practice, we evaluate Eq. (2) using pinhole samples with two spin-up and two spin-down valence neutrons, which constitute the bulk of the pinhole ensemble. Motivated by the dineutron correlations identified in 7H and 8He, we characterize representative four-neutron configurations in terms of two opposite-spin pair coordinates, which are sensitive to dineutron-like structures. Fig. 3 summarizes the resulting four-neutron geometries. We first present in Fig. 3(a) the marginal distribution of inter-pair opening and torsion angles (, ), obtained by integrating the four-body correlation function defined in Eq. (2) over the intra-pair angles (, ) and (, ). Guided by the correlations of 7H discussed above, we identify dineutron-like configurations by requiring pairs to satisfy and when constructing the marginal distribution of the inter-pair opening and torsion angles, with . The deviation of from quantifies the degree of radial asymmetry within each pair. The robustness of these angular distributions with respect to variations in the cuts is demonstrated in Ref. [70], where the normalized marginal four-body correlation distribution changes only weakly and remains consistent within uncertainties. The marginal probability density is strongly enhanced for opening angles in the range –, and forms a broad maximum centered around . This distribution is most naturally associated with configurations containing two dineutron-like pairs on opposite sides of the system c.m., with finite torsion indicating non-planarity. While Ref. [100] suggested possible coplanar “great-circle” configurations in 8He based on a pure shell-model occupation, our results for 7H show little strength near , indicating that planar arrangements are disfavored when many-body correlations are included explicitly. These features, together with the enhanced probability at finite torsion angles, suggest that the inter- geometry itself is modified by many-body correlations. Consequently, the four-neutron subsystem favors an extended, twisted configuration that is approximately spatially symmetric about the system c.m. but non-planar. In addition to this dominant extended configuration, tetraneutron-like geometries corresponding to the region of small opening angles, where the four neutrons reside on the same side of the system c.m., occur with lower probability. As shown in Fig. 3(b), this component accounts for about 5% of the sampled configurations shown in panel (b) and is concentrated at in the range –. Its associated torsion angle also lies predominantly in the range –, indicating a strongly non-coplanar four-neutron geometry (large ) compared to the symmetric – configuration. For comparison, the distribution in 8He is shifted to smaller angles and more strongly suppressed at large , indicating a less prominent symmetric - geometry than in 7H under the same selection. This may reflect core-induced medium effects on the four-neutron configuration, in that the different proton background and core structure modify the environment experienced by the four neutrons.
To further elucidate these two geometries, Figs. 3(c–f) display the internal angular and radial correlations for configurations corresponding to the extended, approximately symmetric – geometry at and the compact tetraneutron-like geometry at (see [70] for additional details). For the extended – geometry, Figs. 3(d,f), each dineutron pair retains a small internal opening angle , with its distribution peaked near , and the two pairs exhibit nearly symmetric distributions of the distances (defined as the distances from each pair c.m. to the system c.m.) that extend to about 4–5 fm. In contrast, the tetraneutron-like geometry, Figs. 3(c,e), shows enhanced probability density near characteristic radial-distance pairs fm and fm from the system c.m., indicating two radially asymmetric pairs located on the same side of the system c.m. Their internal opening-angle distributions are shifted to larger values, with a pronounced maximum near and lying predominantly in the range -. These configurations are also accompanied by large torsion angles with between and . Together, these features lead to a non-coplanar, interlaced geometry. This geometry represents a compact, strongly correlated four-neutron configuration, in which the distinction between the two clusters becomes less pronounced, suggesting enhanced four-body correlations among the four neutrons. It may be viewed as a structural precursor to a possible tetraneutron configuration. Overall, the contrast between these configurations highlights two representative geometries of the correlated four-neutron subsystem in 7H: a spatially symmetric, extended – geometry with large inter-pair separation and finite torsion, which accounts for roughly 95% of the total marginal distribution, and a compact, twisted tetraneutron-like geometry governed by pronounced four-body correlations, which contributes the remaining 5%. The strong dominance of the extended – geometry in the neutron-rich light nuclei studied here implies that experimental searches for tetraneutron-like configurations are likely to receive sizable contributions from symmetric – modes, posing a serious challenge to isolating a genuine tetraneutron signal.
Summary: We perform ab initio NLEFT calculations of the neutron-rich nuclei 6,7H and 8He, including full -body correlations and a Bayesian analysis over an ensemble of chiral interactions. We predict the ground-state energies of the extreme hydrogen isotopes 6,7H and obtain the energies of the benchmark nucleus 8He in agreement with available data, while simultaneously uncovering their internal multi-neutron structure. Intrinsic densities and two-neutron correlation functions reveal a compact triton (alpha) cluster in 7H (8He) surrounded by four valence neutrons that form compact dineutron pairs. Building on this, we construct a reduced four-body correlation function to resolve the geometry of the four-neutron subsystem. We extract the corresponding spatial and angular correlation patterns between each nucleon in two representative geometries. We find that the 7H ground state is dominated (about 95%) by a spatially symmetric, extended – geometry with large inter-pair separation and finite torsion, while a compact, twisted tetraneutron-like configuration with pronounced four-body correlations accounts for the remaining 5%. These results provide a microscopic picture of emergent multi-neutron correlations in 7H and 8He and imply that experimental searches for tetraneutron-like configurations are likely to be contaminated by sizable backgrounds from symmetric – configurations.
Acknowledgements.
We are grateful for discussions with Y. Z. Ma, S. H. Shen, Z. X. Ren, F. Hildenbrand, T. Lähde, A. Sarkar, and other members of the NLEFT collaboration. S. Zhang thanks the discussion with S. W. Huang, S. M. Wang, B. S. Hu, J. G. Li, and X. X. Sun. This work is part of the EXOTIC grant and was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101018170). The work of UGM was also supported by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK project no. 123F464). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Furthermore, the authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG).References
- [1] F. M. Marqués and J. Carbonell, Eur. Phys. J. A 57 (2021) no.3, 105 doi:10.1140/epja/s10050-021-00417-8 [arXiv:2102.10879 [nucl-ex]].
- [2] F. M. Marques, M. Labiche, N. A. Orr, J. C. Angelique, L. Axelsson, B. Benoit, U. C. Bergmann, M. J. G. Borge, W. N. Catford and S. P. G. Chappell, et al. Phys. Rev. C 65 (2002), 044006 doi:10.1103/PhysRevC.65.044006 [arXiv:nucl-ex/0111001 [nucl-ex]].
- [3] K. Kisamori, S. Shimoura, H. Miya, S. Michimasa, S. Ota, M. Assie, H. Baba, T. Baba, D. Beaumel and M. Dozono, et al. Phys. Rev. Lett. 116 (2016) no.5, 052501 doi:10.1103/PhysRevLett.116.052501
- [4] M. Duer, T. Aumann, R. Gernhäuser, V. Panin, S. Paschalis, D. M. Rossi, N. L. Achouri, D. Ahn, H. Baba and C. A. Bertulani, et al. Nature 606 (2022) no.7915, 678-682 doi:10.1038/s41586-022-04827-6
- [5] S. W. Huang, Z. H. Yang, F. M. Marqués, N. L. Achouri, D. S. Ahn, T. Aumann, H. Baba, D. Beaumel, M. Böhmer and K. Boretzky, et al. Few Body Syst. 62 (2021) no.4, 102 doi:10.1007/s00601-021-01691-4
- [6] M. Oertel, M. Hempel, T. Klähn and S. Typel, Rev. Mod. Phys. 89 (2017) no.1, 015007 doi:10.1103/RevModPhys.89.015007 [arXiv:1610.03361 [astro-ph.HE]].
- [7] Y. L. Ye, X. F. Yang, H. Sakurai and B. S. Hu, Nature Rev. Phys. 7 (2025) no.1, 21-37 doi:10.1038/s42254-024-00782-5
- [8] O. Hen, G. A. Miller, E. Piasetzky and L. B. Weinstein, Rev. Mod. Phys. 89 (2017) no.4, 045002 doi:10.1103/RevModPhys.89.045002 [arXiv:1611.09748 [nucl-ex]].
- [9] J. Arrington, N. Fomin and A. Schmidt, Ann. Rev. Nucl. Part. Sci. 72 (2022), 307-337 doi:10.1146/annurev-nucl-102020-022253 [arXiv:2203.02608 [nucl-ex]].
- [10] I. Tanihata, H. Hamagaki, O. Hashimoto, Y. Shida, N. Yoshikawa, K. Sugimoto, O. Yamakawa, T. Kobayashi and N. Takahashi, Phys. Rev. Lett. 55 (1985), 2676-2679 doi:10.1103/PhysRevLett.55.2676
- [11] T. Nakamura, A. M. Vinodkumar, T. Sugimoto, N. Aoi, H. Baba, D. Bazin, N. Fukuda, T. Gomi, H. Hasegawa and N. Imai, et al. Phys. Rev. Lett. 96 (2006), 252502 doi:10.1103/PhysRevLett.96.252502
- [12] I. Tanihata, H. Savajols and R. Kanungo, Prog. Part. Nucl. Phys. 68 (2013), 215-313 doi:10.1016/j.ppnp.2012.07.001
- [13] Y. Kikuchi, K. Ogata, Y. Kubota, M. Sasano and T. Uesaka, PTEP 2016 (2016) no.10, 103D03 doi:10.1093/ptep/ptw148 [arXiv:1603.03858 [nucl-th]].
- [14] A. Revel et al. [R3B], Phys. Rev. Lett. 120 (2018) no.15, 152504 doi:10.1103/PhysRevLett.120.152504 [arXiv:1803.04777 [nucl-ex]].
- [15] Y. Kubota, A. Corsi, G. Authelet, H. Baba, C. Caesar, D. Calvet, A. Delbart, M. Dozono, J. Feng and F. Flavigny, et al. Phys. Rev. Lett. 125 (2020), 252501 doi:10.1103/PhysRevLett.125.252501 [arXiv:2010.04802 [nucl-ex]].
- [16] A. Corsi, Y. Kubota, J. Casal, M. Gómez-Ramos, A. M. Moro, G. Authelet, H. Baba, C. Caesar, D. Calvet and A. Delbart, et al. Phys. Lett. B 840 (2023), 137875 doi:10.1016/j.physletb.2023.137875 [arXiv:2307.06083 [nucl-ex]].
- [17] B. Monteagudo, F. M. Marqués, J. Gibelin, N. A. Orr, A. Corsi, Y. Kubota, J. Casal, J. Gómez-Camacho, G. Authelet and H. Baba, et al. Phys. Rev. Lett. 132 (2024) no.8, 082501 doi:10.1103/PhysRevLett.132.082501 [arXiv:2401.16817 [nucl-ex]].
- [18] M. Hammen, W. Nörtershäuser, D. L. Balabanski, M. L. Bissell, K. Blaum, I. Budinčević, B. Cheal, K. T. Flanagan, N. Frömmgen and G. Georgiev, et al. Phys. Rev. Lett. 121 (2018) no.10, 102501 doi:10.1103/PhysRevLett.121.102501
- [19] S. Geldhof, M. Kortelainen, O. Beliuskina, P. Campbell, L. Cáceres, L. Cañete, B. Cheal, K. Chrysalidis, C. S. Devlin, R. P. de Groote, A. de Roubin, T. Eronen, Z. Ge, W. Gins, A. Koszorus, S. Kujanpää, D. Nesterenko, A. Ortiz-Cortes, I. Pohjalainen, I. D. Moore, A. Raggio, M. Reponen, J. Romero and F. Sommer, Phys. Rev. Lett. 128 (2022) no.15, 152501 doi:10.1103/PhysRevLett.128.152501
- [20] A. J. Miller, K. Minamisono, A. Klose et al., Proton superfluidity and charge radii in proton-rich calcium isotopes, Nat. Phys. 15, 432–436 (2019).
- [21] J. Kahlbow et al. [SAMURAI21-NeuLAND], Phys. Rev. Lett. 133 (2024) no.8, 082501 doi:10.1103/PhysRevLett.133.082501 [arXiv:2407.19303 [nucl-ex]].
- [22] B. Blank and M. Ploszajczak, Rept. Prog. Phys. 71 (2008), 046301 doi:10.1088/0034-4885/71/4/046301 [arXiv:0709.3797 [nucl-ex]].
- [23] M. Pfützner, I. Mukha and S. M. Wang, Prog. Part. Nucl. Phys. 132 (2023), 104050 doi:10.1016/j.ppnp.2023.104050 [arXiv:2304.13391 [nucl-ex]].
- [24] M. Pfutzner, M. Karny, L. V. Grigorenko and K. Riisager, Rev. Mod. Phys. 84 (2012), 567 doi:10.1103/RevModPhys.84.567 [arXiv:1111.0482 [nucl-ex]].
- [25] A. Spyrou, Z. Kohley, T. Baumann, D. Bazin, B. A. Brown, G. Christian, P. A. DeYoung, J. E. Finck, N. Frank and E. Lunderberg, et al. Phys. Rev. Lett. 108 (2012), 102501 doi:10.1103/PhysRevLett.108.102501
- [26] Z. Kohley, T. Baumann, D. Bazin, G. Christian, P. A. DeYoung, J. E. Finck, N. Frank, M. Jones, E. Lunderberg and B. Luther, et al. Phys. Rev. Lett. 110 (2013) no.15, 152501 doi:10.1103/PhysRevLett.110.152501 [arXiv:1303.2617 [nucl-ex]].
- [27] J. Tanaka, Z. H. Yang, S. Typel, S. Adachi, S. W. Bai, P. van Beek, D. Beaumel, Y. Fujikawa, J. X. Han and S. Heil, et al. Science 371 (2021) no.6526, 260-264 doi:10.1126/science.abe4688
- [28] M. Freer, H. Horiuchi, Y. Kanada-En’yo, D. Lee and U.-G. Meißner, Rev. Mod. Phys. 90 (2018) no.3, 035004 doi:10.1103/RevModPhys.90.035004 [arXiv:1705.06192 [nucl-th]].
- [29] Y. Kondo, N. L. Achouri, H. Al Falou, L. Atar, T. Aumann, H. Baba, K. Boretzky, C. Caesar, D. Calvet, H. Chae et al., Nature 620 (2023) no.7976, 965–970 doi:10.1038/s41586-023-06352-6
- [30] A. A. Bezbakh, V. Chudoba, S. A. Krupko, S. G. Belogurov, D. Biare, A. S. Fomichev, E. M. Gazeeva, A. V. Gorshkov, L. V. Grigorenko and G. Kaminski, et al. Phys. Rev. Lett. 124 (2020) no.2, 022502 doi:10.1103/PhysRevLett.124.022502 [arXiv:1906.07818 [nucl-ex]].
- [31] I. A. Muzalevskii, A. A. Bezbakh, E. Y. Nikolskii, V. Chudoba, S. A. Krupko, S. G. Belogurov, D. Biare, A. S. Fomichev, E. M. Gazeeva and A. V. Gorshkov, et al. Phys. Rev. C 103 (2021) no.4, 044313 doi:10.1103/PhysRevC.103.044313 [arXiv:2010.09655 [nucl-ex]].
- [32] T. Shao et al. [A1], Phys. Rev. Lett. 134 (2025) no.16, 162501 doi:10.1103/PhysRevLett.134.162501 [arXiv:2501.01232 [nucl-ex]].
- [33] L. V. Grigorenko and M. V. Zhukov, Phys. Rev. C 68 (2003), 054005 doi:10.1103/PhysRevC.68.054005
- [34] K. Hagino and H. Sagawa, Phys. Rev. C 72 (2005), 044321 doi:10.1103/PhysRevC.72.044321 [arXiv:nucl-th/0508058 [nucl-th]].
- [35] K. Hagino, H. Sagawa, J. Carbonell and P. Schuck, Phys. Rev. Lett. 99 (2007), 022506 doi:10.1103/PhysRevLett.99.022506 [arXiv:nucl-th/0611064 [nucl-th]].
- [36] G. Papadimitriou, A. T. Kruppa, N. Michel, W. Nazarewicz, M. Ploszajczak and J. Rotureau, Phys. Rev. C 84 (2011), 051304 doi:10.1103/PhysRevC.84.051304 [arXiv:1109.0223 [nucl-th]].
- [37] L. V. Grigorenko, I. G. Mukha and M. V. Zhukov, Phys. Rev. Lett. 111 (2013) no.4, 042501 doi:10.1103/PhysRevLett.111.042501 [arXiv:1304.4901 [nucl-th]].
- [38] A. T. Kruppa, G. Papadimitriou, W. Nazarewicz and N. Michel, Phys. Rev. C 89 (2014) no.1, 014330 doi:10.1103/PhysRevC.89.014330 [arXiv:1310.7519 [nucl-th]].
- [39] C. Romero-Redondo, S. Quaglioni, P. Navratil and G. Hupin, Phys. Rev. Lett. 117 (2016) no.22, 222501 doi:10.1103/PhysRevLett.117.222501 [arXiv:1606.00066 [nucl-th]].
- [40] S. M. Wang, N. Michel, W. Nazarewicz and F. R. Xu, Phys. Rev. C 96 (2017) no.4, 044307 doi:10.1103/PhysRevC.96.044307 [arXiv:1707.08954 [nucl-th]].
- [41] Y. Z. Ma, F. R. Xu, N. Michel, S. Zhang, J. G. Li, B. S. Hu, L. Coraggio, N. Itaco and A. Gargano, Phys. Lett. B 808 (2020), 135673 doi:10.1016/j.physletb.2020.135673 [arXiv:2008.01420 [nucl-th]].
- [42] C. A. Bertulani, M. S. Hussein and G. Verde, Phys. Lett. B 666 (2008), 86-90 doi:10.1016/j.physletb.2008.06.062 [arXiv:0801.0784 [nucl-th]].
- [43] D. S. Delion, R. J. Liotta and R. Wyss, Phys. Rev. C 87 (2013) no.3, 034328 doi:10.1103/PhysRevC.87.034328
- [44] T. Oishi, K. Hagino and H. Sagawa, Phys. Rev. C 90 (2014) no.3, 034303 doi:10.1103/PhysRevC.90.034303 [arXiv:1404.3019 [nucl-th]].
- [45] T. Oishi, M. Kortelainen and A. Pastore, Phys. Rev. C 96 (2017) no.4, 044327 doi:10.1103/PhysRevC.96.044327 [arXiv:1606.03111 [nucl-th]].
- [46] S. M. Wang and W. Nazarewicz, Phys. Rev. Lett. 126 (2021) no.14, 142501 doi:10.1103/PhysRevLett.126.142501 [arXiv:2104.03195 [nucl-th]].
- [47] K. Hagino, N. Takahashi and H. Sagawa, Phys. Rev. C 77 (2008), 054317 doi:10.1103/PhysRevC.77.054317 [arXiv:0803.3258 [nucl-th]].
- [48] E. Garrido and A. S. Jensen, J. Phys. G 52 (2025) no.10, 105101 doi:10.1088/1361-6471/ae0959 [arXiv:2509.17908 [nucl-th]].
- [49] Y. Kanada-En’yo, Phys. Rev. C 76 (2007), 044323 [erratum: Phys. Rev. C 101 (2020) no.6, 069902] doi:10.1103/PhysRevC.76.044323 [arXiv:0707.2120 [nucl-th]].
- [50] S. Aoyama and N. Itagaki, Phys. Rev. C 80 (2009), 021304 doi:10.1103/PhysRevC.80.021304
- [51] N. Itagaki, M. Ito, K. Arai, S. Aoyama and T. Kokalova, Phys. Rev. C 78 (2008), 017306 doi:10.1103/PhysRevC.78.017306
- [52] F. Kobayashi and Y. Kanada-En’yo, Phys. Rev. C 88 (2013) no.3, 034321 doi:10.1103/PhysRevC.88.034321 [arXiv:1305.4323 [nucl-th]].
- [53] Y. Yamaguchi, W. Horiuchi, T. Ichikawa and N. Itagaki, Phys. Rev. C 108 (2023) no.1, L011304 doi:10.1103/PhysRevC.108.L011304 [arXiv:2305.15624 [nucl-th]].
- [54] K. Nakagawa and Y. Kanada-En’yo, Phys. Rev. C 112 (2025) no.3, 034309 doi:10.1103/tjnl-jnxp [arXiv:2505.16235 [nucl-th]].
- [55] R. Weiss and S. Gandolfi, Phys. Rev. C 108 (2023) no.2, L021301 doi:10.1103/PhysRevC.108.L021301 [arXiv:2301.09605 [nucl-th]].
- [56] G. Jacob and T. a. j. Maris, Rev. Mod. Phys. 38 (1966), 121-142 doi:10.1103/RevModPhys.38.121
- [57] G. Jacob and T. A. J. Maris, Rev. Mod. Phys. 45 (1973), 6-21 doi:10.1103/RevModPhys.45.6
- [58] K. Yoshida and J. Tanaka, Prog. Theor. Exp. Phys. 2025, ptaf119 (2025) doi:10.1093/ptep/ptaf119 [arXiv:2412.16649 [nucl-th]].
- [59] T. Faestermann, A. Bergmaier, R. Gernhäuser, D. Koll and M. Mahgoub, Phys. Lett. B 824 (2022), 136799 doi:10.1016/j.physletb.2021.136799
- [60] K. Miki et al. [RIBF-SHARAQ11 and RCNP-E502], Phys. Rev. Lett. 133 (2024) no.1, 012501 doi:10.1103/PhysRevLett.133.012501
- [61] R. Lazauskas, E. Hiyama and J. Carbonell, Phys. Rev. Lett. 130 (2023) no.10, 102501 doi:10.1103/PhysRevLett.130.102501 [arXiv:2207.07575 [nucl-th]].
- [62] I. A. Mazur, M. K. Efimenko, A. I. Mazur, I. J. Shin, V. A. Kulikov, A. M. Shirokov and J. P. Vary, Phys. Rev. C 110 (2024) no.1, 014004 doi:10.1103/PhysRevC.110.014004 [arXiv:2403.18232 [nucl-th]].
- [63] T. A. Lähde and U.-G. Meißner, Lect. Notes Phys. 957 (2019), 1-396 Springer, 2019, ISBN 978-3-030-14187-5, 978-3-030-14189-9 doi:10.1007/978-3-030-14189-9
- [64] B. N. Lu, N. Li, S. Elhatisari, D. Lee, E. Epelbaum and U.-G. Meißner, Phys. Lett. B 797 (2019), 134863 doi:10.1016/j.physletb.2019.134863 [arXiv:1812.10928 [nucl-th]].
- [65] B. N. Lu, N. Li, S. Elhatisari, Y. Z. Ma, D. Lee and U.-G. Meißner, Phys. Rev. Lett. 128 (2022) no.24, 242501 doi:10.1103/PhysRevLett.128.242501 [arXiv:2111.14191 [nucl-th]].
- [66] S. Elhatisari, L. Bovermann, Y. Z. Ma, E. Epelbaum, D. Frame, F. Hildenbrand, M. Kim, Y. Kim, H. Krebs and T. A. Lähde, et al. Nature 630 (2024) no.8015, 59-63 doi:10.1038/s41586-024-07422-z [arXiv:2210.17488 [nucl-th]].
- [67] I. Vernon, M. Goldstein and R. G. Bower, Bayesian Anal. 5, no.4, 619–669 (2010) doi:10.1214/10-BA524.
- [68] I. Vernon, M. Goldstein and R. Bower, Statist. Sci. 29, 81–90 (2014) doi:10.1214/13-STS440.
- [69] B. S. Hu, W. G. Jiang, T. Miyagi, Z. H. Sun, A. Ekström, C. Forssén, G. Hagen, J. D. Holt, T. Papenbrock and S. R. Stroberg, et al. Nature Phys. 18 (2022) no.10, 1196-1200 doi:10.1038/s41567-023-02324-9 [arXiv:2112.01125 [nucl-th]].
[…]
- [71] S. Elhatisari, E. Epelbaum, H. Krebs, T. A. Lähde, D. Lee, N. Li, B. n. Lu, U.-G. Meißner and G. Rupak, Phys. Rev. Lett. 119 (2017) no.22, 222505 doi:10.1103/PhysRevLett.119.222505 [arXiv:1702.05177 [nucl-th]].
- [72] R. B. Wiringa, S. C. Pieper, J. Carlson and V. R. Pandharipande, Phys. Rev. C 62 (2000), 014001 doi:10.1103/PhysRevC.62.014001 [arXiv:nucl-th/0002022 [nucl-th]].
- [73] T. Otsuka, T. Abe, T. Yoshida, Y. Tsunoda, N. Shimizu, N. Itagaki, Y. Utsuno, J. Vary, P. Maris and H. Ueno, Nature Commun. 13 (2022) no.1, 2234 doi:10.1038/s41467-022-29582-0
- [74] S. H. Shen, S. Elhatisari, T. A. Lähde, D. Lee, B. N. Lu and U.-G. Meißner, Nature Commun. 14 (2023) no.1, 2777 doi:10.1038/s41467-023-38391-y [arXiv:2202.13596 [nucl-th]].
- [75] S. H. Shen, S. Elhatisari, D. Lee, U.-G. Meißner and Z. X. Ren, Phys. Rev. Lett. 134 (2025) no.16, 162503 doi:10.1103/PhysRevLett.134.162503 [arXiv:2411.14935 [nucl-th]].
- [76] G. F. Bertsch and H. Esbensen, Annals Phys. 209 (1991), 327 doi:10.1016/0003-4916(91)90033-5
- [77] D. V. Aleksandrov, E. A. Ganza, Yu. A. Glukhov, B. G. Novatsky, A. A. Ogloblin, and D. N. Stepanov, Sov. J. Nucl. Phys. 39, 323 (1984) [Yad. Fiz. 39, 513 (1984)].
- [78] A. V. Belozyorov et al., Nucl. Phys. A 460, 352 (1986).
- [79] Y. B. Gurov, D. V. Aleshkin, S. V. Lapushkin, P. V. Morokhov, A. V. Panin, V. A. Pechkurov, N. O. Poroshin, V. G. Sandukovsky, M. V. Telkushev and B. A. Chernyshev, JETP Lett. 78 (2003), 183-187 doi:10.1134/1.1622028
- [80] M. Caamano, D. Cortina-Gil, W. Mittig, H. Savajols, M. Chartier, C. E. Demonchy, B. Fernandez, M. B. G. Hornillos, A. Gillibert and B. Jurado, et al. Phys. Rev. C 78 (2008), 044001 doi:10.1103/PhysRevC.78.044001
- [81] A. A. Korsheninnikov et al., Phys. Scr. T88, 199 (2000).
- [82] M. Caamaño, D. Cortina-Gil, W. Mittig, H. Savajols, M. Chartier, C. E. Demonchy, B. Fernández, M. B. Gómez Hornillos, A. Gillibert, B. Jurado, O. Kiselev, R. Lemmon, A. Obertelli, F. Rejmund, M. Rejmund, P. Roussel-Chomaz and R. Wolski, Phys. Rev. Lett. 99 (2007), 062502 doi:10.1103/PhysRevLett.99.062502
- [83] E. Y. Nikolskii, A. A. Korsheninnikov, H. Otsu, H. Suzuki, K. Yoneda, H. Baba, K. Yamada, Y. Kondo, N. Aoi and A. S. Denikin, et al. Phys. Rev. C 81 (2010), 064606 doi:10.1103/PhysRevC.81.064606
- [84] M. Caamaño, T. Roger, A. M. Moro, G. F. Grinyer, J. Pancin, et al., EPJ Web Conf. 232 (2020), 04002 doi:10.1051/epjconf/202023204002
- [85] Y. H. Lin, H. W. Hammer and U.-G. Meißner, Phys. Rev. Lett. 128 (2022) no.5, 052002 doi:10.1103/PhysRevLett.128.052002 [arXiv:2109.12961 [hep-ph]].
- [86] S. Zhang, S. Elhatisari, U.-G. Meißner and S. H. Shen, Phys. Lett. B 869 (2025), 139839 doi:10.1016/j.physletb.2025.139839 [arXiv:2411.17462 [nucl-th]].
- [87] R. C. Johnson, Phys. Lett. B 114 (1982), 147-151 doi:10.1016/0370-2693(82)90134-4
- [88] B. N. Lu, T. A. Lähde, D. Lee and U. G. Meißner, Phys. Rev. D 90 (2014) no.3, 034507 doi:10.1103/PhysRevD.90.034507 [arXiv:1403.8056 [nucl-th]].
- [89] D. R. Tilley, J. H. Kelley, J. L. Godwin, D. J. Millener, J. E. Purcell, C. G. Sheu and H. R. Weller, Nucl. Phys. A 745, 155-362 (2004) doi:10.1016/j.nuclphysa.2004.09.059
- [90] Z. H. Yang, Y. L. Ye, B. Zhou, H. Baba, R. J. Chen, Y. C. Ge, B. S. Hu, H. Hua, D. X. Jiang and M. Kimura, et al. Phys. Rev. Lett. 131 (2023) no.24, 242501 doi:10.1103/PhysRevLett.131.242501
- [91] N. Keeley, F. Skaza, V. Lapoux, N. Alamanos, F. Auger, A. Drouart, A. Gillibert, L. Nalpas, E. C. Pollacco and R. Raabe, et al. Phys. Lett. B 646 (2007), 222-226 doi:10.1016/j.physletb.2007.01.035
- [92] F. Skaza, N. Alamanos, F. Auger, A. Drouart, A. Gillibert, V. Lapoux, L. Nalpas, E. Pollacco, N. Keeley and R. Wolski, et al. Nucl. Phys. A 788 (2007), 260-265 doi:10.1016/j.nuclphysa.2007.01.093
- [93] M. Holl, R. Kanungo, Z. H. Sun, G. Hagen, J. A. Lay, A. M. Moro, P. Navrátil, T. Papenbrock, M. Alcorta and D. Connolly, et al. Phys. Lett. B 822 (2021), 136710 doi:10.1016/j.physletb.2021.136710 [arXiv:2110.01592 [nucl-ex]].
- [94] K. Fossez, J. Rotureau and W. Nazarewicz, Phys. Rev. C 98 (2018) no.6, 061302 doi:10.1103/PhysRevC.98.061302 [arXiv:1806.02936 [nucl-th]].
- [95] H. H. Li, J. G. Li, N. Michel and W. Zuo, Phys. Rev. C 104 (2021) no.6, L061306 doi:10.1103/PhysRevC.104.L061306 [arXiv:2305.05414 [nucl-th]].
- [96] E. Hiyama, R. Lazauskas and J. Carbonell, Phys. Lett. B 833 (2022), 137367 doi:10.1016/j.physletb.2022.137367 [arXiv:2207.04634 [nucl-th]].
- [97] M. Assie, J. A. Scarpaci, D. Lacroix, J. C. Angelique, D. Bazin, D. Beaumel, Y. Blumenfeld, W. N. Catford, M. Chabot and A. Chatterjee, et al. Mod. Phys. Lett. A 25 (2010), 1846-1849 doi:10.1142/S0217732310000460
- [98] P. Mueller, I. A. Sulai, A. C. C. Villari, J. A. Alcantara-Nunez, R. Alves-Conde, K. Bailey, G. W. F. Drake, M. Dubois, C. Eleon and G. Gaubert, et al. Phys. Rev. Lett. 99 (2007), 252501 doi:10.1103/PhysRevLett.99.252501 [arXiv:0801.0601 [nucl-ex]].
- [99] M. V. Zhukov, A. A. Korsheninnikov and M. H. Smedberg, Phys. Rev. C 50 (1994), R1-R4 doi:10.1103/PhysRevC.50.R1
- [100] P. Mei and P. Van Isacker, Annals Phys. 327 (2012), 1182–1201 doi:10.1016/j.aop.2011.12.003.
I Supplemental Material
I.1 Uncertainty quantification of chiral forces
We start from 284 non-implausible N3LO chiral interactions on the lattice [1], obtained via history matching [2, 3, 4, 1] that explicitly account for uncertainties in NN scattering data, the truncation of the chiral hierarchy, and binding energies of selected nuclei. To construct the posterior distribution informed by experimental data, the first step typically involves incorporating additional observables associated with the relevant structural features, thereby conditionally refining the physical credibility of the non-implausible interaction set. The posterior is subsequently obtained within a Bayesian framework that combines a Gaussian likelihood with a prior over the chiral interactions,
where denotes a specific chiral interaction from the non-implausible ensemble, and are the calculated and experimental observables, and and represent their theoretical and experimental uncertainties, respectively. The prior , derived from the values of the history-matching fits [1], reflects the relative credibility of each interaction in reproducing the calibration data, while the likelihood quantifies the agreement between the predictions of each interaction and the selected observables.
In this work, because the 3H binding energy relevant to the hydrogen isotopes was already used in the history-matching procedure that constrained the chiral interactions, the likelihood is approximated as uniform. For each chiral interaction, we calculate the ground-state energies of 6,7H and extract their extrapolated values with statistical uncertainties. The total uncertainty originates from several sources, including the many-body method approximation (first-order perturbative treatment in the wavefunction matching), finite-volume extrapolation, Euclidean-time projection, and statistical noise. The systematic uncertainty associated with the many-body method approximation is reduced through variational Hamiltonian optimization, resulting in a residual error smaller than the statistical one [1]. As demonstrated below, finite-volume effects also lie within the statistical uncertainties. Therefore, we adopt the statistical uncertainty as the representative error throughout this work. Based on the posterior distribution , we generate 80000 random samples by sampling Gaussian fluctuations around the extrapolated values according to their standard deviations, yielding the final statistical distributions and correlations among the observables.
I.2 Evaluation of observables with pinhole configurations
In the pinhole method, we first introduce the normal-ordered -body density operator , where is the one-body density operator for nucleons with spin and isospin at lattice site in coordinate space. A key feature of the pinhole method is that the exact locations of nucleons with spin and isospin can be sufficiently sampled by inserting the normal-ordered -body density operator. Then, we insert the operator at middle time of the Euclidean time projection amplitudes,
where are the initial and final trial wave functions, propagated via the transfer matrix . Here, denotes the simple Hamiltonian used in the propagation, while denotes the unitarily transformed full Hamiltonian obtained by wave-function matching, as discussed in Ref. [1]. The transfer matrix is connected with the evolution operator via Trotterization, which approximates the time evolution by dividing it into discrete steps for computational feasibility. Here, the difference is calculated by perturbation theory, see Ref. [5] for more details. In the large limit, we can obtain the required physical states. Due to the completeness relation of the -body density operator , the sum of this expectation value over and yields times the amplitude , which can be calculated using Auxiliary Fields Monte Carlo simulations. In the auxiliary field formalism, the transfer matrix at each time slice depends on the auxiliary fields , and Eq. (S2) can be rewritten in terms of the auxiliary fields. To evaluate it, we perform importance sampling according to the amplitude,
which reflects the -body density information with the auxiliary fields , pinhole locations and spin-isospins indices . To generate different amplitudes, the auxiliary fields are updated by the shuttle algorithm [5], and then the pinhole locations and spin-isospins indices are sampled by the Metropolis algorithm. For any operator that does not mix pinhole configurations, we evaluate its expectation value as
where the summation symbol indicates the sum of different auxiliary fields , pinhole locations and spin-isospins indices , and is the complex phase of each pinhole amplitude.
I.3 Intrinsic densities
Recent experimental observations have provided growing evidence for intrinsic clustering in nuclei [6]. On the theoretical side, many-body approaches based on core or reference-state frameworks have achieved remarkable success in describing nuclear structure [7, 8, 9, 10, 11]. For the light nuclei studied in this work, it is therefore of particular interest to investigate the possible presence of preformed clusters and their role in shaping the intrinsic configurations of these systems. For each sampled pinhole configuration, we construct all nucleon triplets or quadruplets that satisfy the spin–isospin coupling required for the cluster. Since the clusters we aim to identify are expected to be relatively compact and approximately Gaussian in coordinate space [12, 13, 14], we assign to each candidate cluster a Gaussian-type probability weight based on the distances of its nucleons from the cluster center of mass (c.m.)
where denotes the position of the th nucleon in the cluster and is the c.m. coordinate of the cluster. In the present work, the candidate with the largest probability weight is selected as the intrinsic cluster, while more sophisticated selection schemes can be explored in future work.
I.4 Schematic illustration of four-neutron correlations in 7H
Figure S1 schematically illustrates the angular coordinates used to characterize the four-body correlations in light nuclei, using 7H as an example. The four neutrons can be partitioned into two neutron-neutron pairs. We adopt the same two-body angular definitions for each pair as in the two-body correlation function, and define the pair vectors and as the relative displacement vectors connecting the two neutrons within each pair. The intra-pair opening angles and quantify the internal compactness of the pairs, while and encode the degree of radial asymmetry between the two neutrons within a pair (as in the two-body analysis). The angle characterizes the relative opening between the pair c.m. vectors, and the torsion angle measures the non-coplanarity (twist) between the two neutron-neutron pair vectors. This set of angles provides a convenient parameterization of the four-neutron geometry and underlies the angular distributions discussed in the main text.
I.5 Extrapolation in Euclidean time
Since it is computationally infeasible to evolve the Hamiltonian to infinite Euclidean time, an extrapolation to the limit is required. Based on spectral decomposition of energy at Euclidean time [15, 16], the expectation value of the Hamiltonian can be written as
where denotes the -th eigenstate, is its energy gap to the g.s., and represents its overlap with the initial state relative to that of the g.s.. Following the procedure in Ref. [17], we first fit the energies calculated with the simple Hamiltonian using the lowest order of Eq. (S6) to extract and . These parameters are then used in a second-stage fit for general operators, such as higher-order energy correlations and observables that commute with the Hamiltonian, based on the functional form
I.6 Finite-volume effects
To assess finite-volume effects, we computed the ground-state energy of 7H on three lattice sizes , , and , using two fixed Euclidean times chosen to balance physical reliability and statistical uncertainty (see Table S1). The ground-state energies show a mild dependence on the lattice size. Specifically, the results differ slightly from those at by less than 0.05 MeV for both MeV-1 and MeV-1, and differ somewhat more from by about 0.08 MeV, all within statistical uncertainties. These findings indicate that finite-volume effects at are minor, and no volume extrapolation is performed in this work. A more systematic treatment of finite-volume effects will be required in future calculations with higher statistical precision.
| [fm] | (MeV-1)[MeV] | (MeV-1)[MeV] |
| 14.52 | ||
| 15.84 | ||
| 18.48 |
I.7 Experimental information on 6H and 7H
Table S2 summarizes experimental determinations of the energies and widths of 6H and 7H obtained with different reaction mechanisms. Energies and widths are given relative to the 3H+ threshold for 6H and the 3H+ threshold for 7H. For 6H and 7H, we list only those measurements that report candidate ground-state energies and, where available, widths. Other studies have provided evidence for the existence of these systems without extracting specific energies [31], and some experiments do not report any clear signal attributable to 6H or 7H under their experimental conditions [32, 33, 34, 35]. The values measured in different reactions illustrate that the experimentally reported ground-state energies and widths of 6H and 7H differ significantly across reaction mechanisms and analyses.
| Nucleus | (MeV) | (MeV) | Exp. |
| 6H | 7Li(7Li,8B)6H [18] | ||
| 6H | 9Be(11B,14O)6H [19] | ||
| 6H | 9Be(,)6H [20] | ||
| 6H | 11B(,4He)6H [20] | ||
| 6H | 12C(8He,14N)6H [21] | ||
| 6H | – | 2H(8He,4He)6H [22] | |
| 6H | 7Li()6H [23] | ||
| 7H | 2.0 | p(8He,2He)7H [24] | |
| 7H | – | – | 2H(8He,7H)3He [25] |
| 7H | 12C(8He,13N)7H [26] | ||
| 7H | – | 2H(8He,3He)7H [27] | |
| 7H | – | 2H(8He,3He)7H [28] | |
| 7H | 12C(8He,13N)7H [29] | ||
| 7H | – | 2H(10Be,3He)7H [30] |
I.8 Two-neutron correlations in 6He and 3H
Figure S2(a) shows the two-neutron correlation density of the valence neutrons in 6He. Since 6He has been extensively studied both experimentally and theoretically [36, 37, 38], it provides a benchmark for our correlation-analysis procedure. Figure S2(b) shows the two-neutron correlation density of 3H for comparison with 7H. Note that the 7H correlations are normalized by the total number of neutron pairs, . Therefore, the per-pair short-distance component shown for 7H is reduced by a factor of 15 relative to the single-pair 3H benchmark purely due to this normalization convention. The two-neutron correlation density characterizes the spatial extent and angular structure of the neutron pair in 3H and exhibits a similar structure to the short-distance component of the multi-neutron correlation patterns obtained in 7H.
I.9 Intrinsic densities, two-neutron correlations, and dineutron sizes in 8He and 7H
Figure S3 summarizes the intrinsic nucleon densities, two-neutron correlations, and resulting dineutron-size distributions for 8He and 7H, analyzed within the same intrinsic-frame definitions as in the main text. Panel (a) shows the intrinsic total-nucleon and proton densities, visualized as slices at and fm. The intrinsic densities indicate a compact -like cluster, surrounded by four valence neutrons, with a more compact valence-neutron distribution than in 7H. Panel (b) displays the two-neutron correlation density for the four valence neutrons in 8He, where denotes the distance of each neutron from the 8He c.m. and is the opening angle between the two neutron positions. When the two-neutron correlations are averaged over all valence-neutron pairs in 8He, the correlation density exhibits a single broad maximum around in our calculation. In the idealized limit of a pure shell-model configuration, this averaging largely washes out orientation dependence, yielding an approximately isotropic neutron-neutron angular density. Consequently, the resulting distribution is dominated by the phase-space factor . The contrast with 7H is instructive: while both systems deviate from the idealized limit, the departure is substantially stronger in 7H, allowing nontrivial angular structures to survive the pair averaging, whereas 8He remains comparatively closer to a dominant configuration. In this analysis, we focus on valence-neutron pairs with opposite spin projections. This selection suppresses the largely uncorrelated background (including cross pairs spanning different correlated clusters) and enhances spatially clustered configurations. As shown in panel (b), this conditional correlation exhibits a characteristic ridge at small opening angles , signaling compact dineutron-like configurations, together with a pronounced enhancement toward large , consistent with the bimodal pattern discussed for 7H. Panel (c) compares the normalized dineutron size distributions for dineutron-like configurations in 8He and 7H selected by combined cuts on the internal opening angle and the angle (see legend for the specific cut choices). Here, we define the dineutron size as the half-separation , whereas the dineutron size often quoted in the literature refers to the full relative distance . From these distributions we extract the root-mean-square dineutron radius, , which is indicated by the vertical lines and quoted in the legend. The extracted shows only a weak dependence on the cuts and remains consistent within uncertainties, indicating that the inferred dineutron size is stable against the selection procedure. This also suggests that our choice of angular windows lies in a reasonable range. Moreover, the extracted dineutron size in 8He corresponds to a typical dineutron size of fm, consistent with the commonly considered range to be fm [39, 40] at the nuclear surface, whereas in 7H it exhibits a larger typical scale of fm.
I.10 Opening angle and dependence in four-neutron correlation of 7H
To assess potential systematic effects associated with the marginal distribution, we examine the stability of the opening-angle distribution with respect to variations of the intra-pair neutron-neutron angle selection and the intra-pair neutron-neutron opening-angle cut. Here, represents the angle between the relative position vector of the two nucleons and the bisector of their opening angle ( quantifies the deviation from radial symmetry). As shown in Fig. S4(a), changing from to at fixed leaves the normalized distribution essentially unchanged. When we vary the selection window, it primarily changes the effective statistics (i.e., the number of accepted configurations) but leaves the relative distribution essentially unchanged, indicating no discernible change in the underlying four-neutron correlation pattern within uncertainties. In this work, we select to balance the radial symmetry and statistics. In Fig. S4(b), we vary the intra-pair neutron-neutron opening-angle cut while keeping fixed. Because directly quantifies the internal geometry of an neutron-neutron pair, this cut defines which intra-pair correlation patterns are included in the sampled configurations. Guided by the two-body correlation analysis discussed above, we find that the dominant strength associated with compact, dineutron-like correlations in 7H is concentrated primarily within , which motivates our baseline choice. Accordingly, for moderate variations of the cut (–), the normalized distribution exhibits only minor differences and remains consistent within uncertainties. This indicates that our baseline selection captures the relevant neutron-neutron correlation physics without either overly truncating compact configurations or introducing spurious distortions.
I.11 Internal angular and radial distributions from the four-neutron correlations in 7H
Figure S5 provides additional information on the internal angular and radial distributions in the four-neutron correlation analysis and serves as a consistency check against the two-neutron correlations in 7H. Panels (a,e) compile the joint probability densities and for the two neutron-neutron pairs identified in each sampled four-neutron configuration. The bulk of the probability density is concentrated around the typical dineutron angular and radial scales extracted independently from the two-neutron correlations (dashed lines), indicating that the leading internal scales inferred from two-body observables are consistent with those obtained from the explicit four-body correlation. This agreement supports the interpretation of the observed bimodal angular pattern in 7H from the two-body correlation analysis: the two-body correlation function represents a statistical distribution over all neutron-neutron pairs and thus provides a probe sensitive to the underlying multi-neutron geometry. Panels (b,f), obtained with the selection , are similar to (a,e) and contain the majority of the sampled configurations, indicating that the (extended) sector dominates the sampled four-body ensemble. Panels (d,h) correspond to a characteristic window around a maxima of the marginal distribution (). For the large- window (d,h), the radial distribution shows that and are comparable, and shifts toward smaller intra-pair opening angles, consistent with a near back-to-back arrangement of the two pair c.m. vectors at comparable radii and a more compact intra-pair geometry. In contrast, the small- window (c,g) shows a more asymmetric joint radial distribution, featuring enhanced weight near fm, alongside a nearly symmetric contribution at fm. The intra-pair opening-angle distribution shifts to larger angles and becomes noticeably broader, with enhanced support around , compared to the tighter concentration near in panel (d). The associated torsion-angle distribution is also weighted toward large in our calculation, indicating a strongly non-coplanar (twisted) and interlaced four-neutron geometry with pronounced four-body correlations.